### v2

# Index Terms—  Visual Document Understanding, Audio Visual Question Answering, Multimodal RAG

Abstract
Document images encapsulate a wealth of knowledge, while the portability of spoken queries enables broader and flexible application scenarios. Yet, no prior work has explored knowledge base question answering over visual document images with queries provided directly in speech. We propose TextlessRAG, the first end-to-end framework for speech-based question answering over large-scale document images. Unlike prior methods, TextlessRAG eliminates ASR, TTS and OCR, directly interpreting speech, retrieving relevant visual knowledge, and generating answers in a fully textless pipeline. To further boost performance, we integrate a layout-aware reranking mechanism to refine retrieval. Experiments demonstrate substantial improvements in both efficiency and accuracy. To advance research in this direction, we also release the first bilingual speech–document RAG dataset, featuring Chinese and English voice queries paired with multimodal document content. Both the dataset and our pipeline will be made available at repository:https://github.com/xiepeijinhit-hue/textlessrag
Index Terms—  Visual Document Understanding, Audio Visual Question Answering, Multimodal RAG
1 Introduction
Fig. 1: A Chinese example of TextlessRAG.
The development of Vision-Language Models (VLMs) has been remarkably rapid[1, 2, 3, 4]. With the introduction of dynamic-resolution visual encoders, recent VLMs have made significant advances in understanding text-rich images[5] These models not only achieve higher accuracy in direct visual document understanding but also eliminate the need for traditional document parsing steps, thereby making the entire process more efficient.
Current research is increasingly shifting towards more complex long-document understanding tasks, including long-form question answering[6, 7] and knowledge base QA from RAG[8, 9, 10]. However, no prior work has explored speech-based querying over document images; existing studies on spoken input have been limited to extending visual question answering on natural images[11, 12]. Spoken knowledge base QA is particularly appealing: it requires neither typing queries on a keyboard nor uploading document images. Instead, users can obtain answers simply by asking questions verbally—an interaction mode that can greatly expand the application scenarios of multimodal large language models.
To this end, we propose TextlessRAG, the first speech-based document visual RAG pipeline. As illustrated in Figure 1, the user poses a question through spoken input. The model then retrieves candidate pages from the knowledge base, followed by layout-aware decomposition of these pages into fine-grained evidence units (e.g., charts, tables, natural images, and textual segments). Based on the selected evidence, the model generates an answer and delivers the response in both speech and text. Notably, the entire pipeline is fully de-textualized—operating without OCR, ASR, or TTS—thereby ensuring both efficiency and robust comprehension of multimodal knowledge, particularly for structured content such as tables and charts.
We further construct the first speech-based visual document RAG benchmark (SV-DOC), which comprises speech-augmented extensions of existing English visual document QA datasets[13, 14, 15, 7, 6] and RAGBench[10], along with a manually annotated Chinese Document RAG dataset (CDR). Experimental results demonstrate that our speech-based visual RAG approach delivers significant improvements in both answer quality and response latency.
The main contributions of this work are summarized as follows:
  • We introduce TextlessRAG, the first end-to-end speech-based visual document RAG pipeline, which operates entirely without ASR, OCR, or TTS.
  • We present SV-DOC, the first benchmark for speech-based visual document RAG, featuring bilingual datasets in English and Chinese.
  • We design a fully text-free workflow augmented with a layout-aware reranking strategy, achieving substantial improvements in both efficiency and accuracy across the RAG process.
2 Method
As shown in Figure 2, TextLessRAG is designed to be both concise and efficient focusing specifically on the retrieval and QA generation components. On Retrieval side, we employ ColQwen-Omnihttps://huggingface.co/vidore/colqwen-omni-v0.1 as a retrieve encoder $\mathrm{Enc}(\cdot)$ to encode both document image pages from document knowedge base $\mathcal{I}=\{P_{1},P_{2}...P_{n}\}$and spoken query $q$ to embeddings as $E=\{e_{1},e_{2}...e_{n}\}$ and $e_{q}$. Then compute the MaxSim Score like ColBerthttps://github.com/stanford-futuredata/ColBERT to rank the top-k candidate pages $T_{k}$:
$E=\{e_{i}=\mathrm{Enc}(I_{i})\mid I_{i}\in\mathcal{I}\},\quad e_{q}=\mathrm{Enc}(q)$ (1)
$s_{i}=Score(P_{i},q)=\mathrm{MaxSim}\big{(}{e_{q}},e_{i}\big{)},\quad e_{i}\in E$ (2)
$\mathcal{T}_{k}=\operatorname{TopK}_{P_{i}\in\mathcal{I}}\Big{(}Score(P_{i},q)\Big{)}=[P_{t_{1}},P_{t_{2}}\dots P_{t_{k}}]$ (3)
On Generation side, we employ Qwen2.5-Omnihttps://huggingface.co/Qwen/Qwen2.5-Omni-7B as a Generator $Gen(\cdot)$, which directly takes the spoken query $q$ together with the top-k re-ranked document images $\mathcal{T}_{k}^{\prime}$ as input, and ultimately produces a spoken answer $\mathcal{A}ns$:
$\mathcal{A}ns=Gen(\mathcal{T}_{k}^{\prime},q),\quad\mathcal{T}_{k}^{\prime}=[P_{t_{1}}^{\prime},P_{t_{2}}^{\prime}\dots P_{t_{k}}^{\prime}]$ (4)
We lay out each top-k page $P_{t}$ into content blocks of chart, table, text and natural image sets by DocLayout-Yolohttps://github.com/opendatalab/DocLayout-YOLO:
$\mathcal{C}ontent=\{\text{chart},\text{table},\text{text},\text{image}\}$ (5)
$YOLO(P_{t})=\bigcup_{c\in\mathcal{C}ontent}Block^{c}(P_{t}),\quad P_{t}\in\mathcal{T}_{k}$ (6)
and refine them with threshold $\theta$ as a lower bound to guarantee the MaxSim Score with query $s_{b}$:
${Block}^{c}_{\theta}=\bigcup_{c\in\mathcal{C}ontent}\left\{b\in Block^{c}(P_{t})\;\middle|\;s_{b}\geq\theta\right\}$ (7)
Then, we resort the image inputs $P_{t}$ to refine representations $P_{t}^{\prime}$ that are more relevant to the given question:
$P_{t}^{\prime}\;=\;\operatorname{Sort}_{\downarrow Score}\big{(}{Block}^{c}_{\theta}\big{)},$ (8)
Overall, the end-to-end process facilitates a fully text-free workflow: it requires neither ASR nor OCR and obviates the need for text-to-answer TTS, enabling direct spoken queries and responses. Furthermore, the layout segmentation and embedding of knowledge base document images are preprocessed in advance. Each query involves two retrieval stages, which not only improves the relevance of retrieved results but also substantially reduces computational overhead.
Fig. 2: Pipeline of TextlessRAG.
3 Data Engine
We developed a data engine to construct the SV-DOC benchmark. As summarized in Table 1, SV-DOC comprises (i) extensions of existing English benchmarks augmented with TTS, and (ii) CDR, the first Chinese Document RAG dataset created from scratch. CDR is built following a five-step pipeline (illustrated on the right of Figure 3), while the final two steps are additionally used to generate speech-augmented English samples.
Specifically, (1) we collect a diverse set of documents and reports across multiple domains and formats, including PDFs and PPTs (see the domain word cloud in Figure 3, left). (2) Each page is segmented into fine-grained units—tables, charts, images, and text blocks—using DocLayout-YOLO. (3) These segments are processed by commercial vision–language agents via batch APIs to generate candidate QA pairs. (4) The QA pairs are refined through rule-based filtering and professional human annotation. (5) Finally, the text QA pairs are converted into speech using Doubao’s TTS API, leveraging over 200 professional voice types.https://www.volcengine.com/docs/6561/1257544.
For the English datasets, we select refined, relevant visual document QA datasets—ChartQA[13], DUDE[15], InfoVQA[14], and SlideVQA[7](following VDocRAG[9]), as well as MMLongBench-Doc (MMLong)[6] and Vidoseek[10] — and apply steps 4 and 5 to augment them with audio queries. Further details of the data construction process and representative examples are provided in the repository.
Fig. 3: Word Cloud of domains of Chinese samples on the left and the 5 steps data engine on the right.
Dataset QA Pool Content Domain
Chartqa[13] 150 119 C Academic
Infovqa[14] 1048 300 C/T/I Infographic
Slidevqa[7] 760 657 C/T/I/X Slide
DUDE[15] 496 422 C/T/X Open
MMLong[6] 1091 5134 C/T/I/X Open
Vidoseek[10] 1142 5349 C/T/I Open
CDR 1260 30583 C/T/I/X Open
Total 5947 42564 C/T/I/X Open
Table 1: Data statistics of SV-Doc Bench.“QA” and “Pool” calculate the amount of qa pairs and images of retrieval pool. The multimodal content of chart, tabel, natural image and text are abbreviated as “C”, “T”, “I” and “X”.
4 Experiment & Result
Model Retriver Query Doc ChartQA DUDE Infovqa SlideVQA MMLong Vidoseek CDR
BM25 - T T 54.8 57.2 50.2 40.7 18.5 84.5 54.9
E5 BERT T T 74.9 40.6 42.5 50.8 23.4 63.5 62.6
NV-Embed-v2 Mistral T T 75.3 43.0 56.5 61.7 38.7 90.3 69.3
CLIP Scratch T I 54.6 23.2 29.7 38.6 17.3 35.8 32.5
DSE Phi3V T I 72.7 55.5 67.4 73.0 43.6 89.4 77.1
VisRAG-Ret MiniCPM-V T I 87.2 56.4 71.9 74.3 53.1 91.2 80.9
VDocRAG Phi3V T I 86.0 57.7 72.9 77.3 49.2 92.8 82.4
ViDoRAG Colqwen2 T I 100 96.5 97.8 96.9 67.0 94.3 87.7
TextLessRAG Colqwen-Omni A I 99.3 91.5 91.6 94.2 66.5 95.4 87.4
Table 2: Rertrive result. “T”, “I”, and “A” represent the query and documents in the form of text, image, and audio, respectively.
Model Generator ChartQA DUDE InfoVQA SlideVQA MMLong Vidoseek CDR
TextRAG Phi3 28.0 40.1 40.5 28.6 6.9 29.8 10.5
TextRAG† 36.6 55.9 45.6 27.8 13.1 31.7 18.7
VDocRAG Phi3V 52.0 48.5 56.2 48.0 14.5 52.1 22.3
VDocRAG† 74.0 66.4 64.6 56.4 21.7 63.8 34.6
ViDoRAG Qwen2.5-VL7B 84.6 86.7 79.1 82.5 37.9 85.7 50.0
ViDoRAG† 84.6 87.4 82.6 84.2 47.3 86.4 70.1
TextLessRAG Qwen-Omni 87.3 78.5 74.5 79.7 33.4 90.2 43.5
TextLessRAG* 87.3 81.3 79.4 82.6 36.7 93.4 47.2
TextLessRAG† 87.3 84.0 80.6 81.8 43.2 92.6 61.3
Table 3: QA result measured by GPT-4o. For each model, we evaluated the QA performance under two input settings: top-5 retrieved images and gold page marked with “†”. Asterisk “*” indicates the top-5 results were further refined by layout reranking. We highlight the best results obtained with gold and top-5 page inputs in yellow and green, respectively.
We evaluate our approach along three dimensions: Retrieval, QA, and Latency. Retrieval is measured using nDCG@5, while QA accuracy is assessed based on GPT-4o’s averaged judgments. Latency is benchmarked on a single 80GB A100 GPU to ensure consistency.
4.1 Retrival Result
Table 2 reports Retrieval results, comparing three text-only methods—BM25[16], E5[17], and NV-Embed-v2[18]— against five visual models(with ground truth text input as query) — CLIP[19], DSE[20], VisRAG-Ret[8], VDocRAG[9], and ViDoRAG[10]—as baselines. For text-based retrieval, OCR text was extracted using Tesseracthttps://github.com/tesseract-ocr/tesseract. Retrieval experiments were then conducted independently within each retrieval pool across seven datasets.
Our proposed method, TextLessRAG, which adopts a novel speech-based retrieval paradigm, demonstrates competitive performance. Compared with three text-to-text retrieval approaches, it consistently and substantially outperforms all of them. This indicates that storing the knowledge base as image embeddings directly encoded by a vision encoder yields stronger representational capacity. Such an approach not only eliminates the need for OCR but also facilitates query retrieval over multimodal content such as tables and charts.
When compared with text-to-image retrieval methods, TextLessRAG—although not surpassing the SOTA ViDoRAG overall—outperforms the other four baselines and achieves the best results on dataset Vidoseek. Moreover, on ChartQA, SlideVQA, MMlong and CDR, the performance gap with SOTA remains within 2%. These findings highlight the significant potential of replacing text with speech for image-based knowledge retrieval. Furthermore, when textual queries are converted into speech, the loss gap from TTS in retrieval accuracy seemed minimal, enabling direct retrieval with queries spoken in diverse voices while maintaining performance comparable to SOTA models.
Based on the model performance consistency across datasets, we observe that our CDR dataset is easier than Vidoseek but more challenging than MMLong. Notably, even with the retrieval pool expanded fivefold(from 5k to 30k), the TextLessRAG retrieval method maintained an accuracy above 87, underscoring its strong capability.
4.2 QA Generation Result
Table 3 presents the end-to-end QA performance results. Although TextLessRAG does not achieve the best performance across all datasets, it substantially outperforms purely text-based methods and even reaches state-of-the-art results on certain datasets, such as 87.3 on ChartQA and 93.4 on Vidoseek. Using golden-label pages as the input, which represent the upper bound of the generator, we find that it is slightly inferior to ViDoRAG on most datasets. Although converting text to speech introduces a minor performance drop, TextLessRAG still substantially outperforms the rest models. With top-5 inputs, it also achieves the best performance on InfoVQA of 79.4 and SlideVQA of 82.6.
We then examine the effect of reranking and observe consistent improvements in the answers across all datasets. Notably, on the InfoVQA and SlideVQA datasets, reranking reverses the advantage of the gold setting, with the performance gains from top-5 inputs leading to superior results. Remarkably, even on the Vidoseek, where retrieval quality is already strong, reranking enables the top-5 answers to surpass those obtained with gold inputs.
These findings indicate that across the entire pipeline, the RAG and QA upper bound performance of TextLessRAG is comparable to that of the SOTA model. This further underscores the great potential of using spoken queries in the QA stage. Moreover, layout reranking, which leverages speech to select image content with finer granularity, enhances both the potential and the overall quality of answers.
4.3 Accuracy and Latency Analysis
Fig. 4: End-to-end accuracy and latency analysis.
Figure4 provides an evaluation of end-to-end RAG in terms of both answer quality and latency. On the left, we jointly evaluate the end-to-end response latency and answer accuracy to assess the overall pipeline. We observe that TextLessRAG strikes a favorable balance between effectiveness and efficiency: compared with ViDoRAG, it achieves a substantial speedup while incurring only a negligible loss in accuracy. VdocRAG is slightly slower than ViDocRAG but exhibits a steep drop in accuracy, while the pure text method is both slower and more error-prone.
In Figure4(right), we break down the latency of each component to identify bottlenecks. Excluding communication overhead, we find that ASR, TTS, and OCR contribute the largest share of latency. By contrast, our textless design leverages direct speech encoding for both retrieval and generation, operating directly on visual features for retrieval and response generation. This end-to-end de-textualized pipeline not only eliminates additional computational and communication costs but also avoids error propagation across cascaded stages, ultimately enabling more accurate and faster responses.
5 Conclusion
In this work, we present TextLessRAG, the first RAG pipeline for visual document knowledge bases with speech-based input and output. We also introduce the first bilingual benchmark for this task and release the first open-source Chinese visual document RAG dataset. Experimental results demonstrate that our textless design eliminates the need for OCR, ASR, and TTS, substantially reducing inference time. Moreover, with the aid of a layout-based reranking strategy, our approach ensures fine-grained block-level awareness in retrieval and generates highly accurate answers. Overall, this pioneering exploration of speech-driven visual RAG reveals the significant potential of speech-based knowledge base question answering and we believe it will facilitate broader and more practical applications of multimodal large language models.